Skip to content

Observatory Status

Siding Spring Observatory - NSW, Australia

COJ

Clamshell
2m
Spectral (fs01)
FLOYDS (en12)
0.4m A
SBIG (kb26)
0.4m B
SBIG (kb83)
Dome A
1m
Sinistro (fa12)
Dome B
1m
Sinistro (fa19)

Wise Observatory - Israel

TLV

Dome A
1m
NRES (nres-agu-ak15)
NRES Igloo

SAAO - Sutherland, South Africa

CPT
SIte-wide power outage since
2023-05-29

Dome A
1m
Sinistro (fa14)
Dome B
1m
Sinistro (fa01)
NRES (nres-agu-ak06)
Dome C
1m
Sinistro (fa06)
NRES (nres-agu-ak05)
Aqawan A
0.4m A
None (sq38)
0.4m B
FLI PL230 (ff13,14,15,16)
ASAS-SN
NRES Igloo

Teide Observatory - Tenerife, Spain

TFN

Dome A
1m
Sinistro (fa20)
Dome B
1m
Sinistro (fa11)
Aqawan A
0.4m A
Not in service since 2023-04-27
SBIG (kb95)
0.4m B
SBIG (kb24)

CTIO - Region IV, Chile

LSC

Dome A
1m
Sinistro (fa15)
Dome B
1m
Sinistro (fa04)
NRES (nres-agu-ak17)
Dome C
1m
Sinistro (fa03)
NRES (nres-agu-ak02)
Aqawan A
0.4m A
SBIG (kb98)
0.4m B
FLI PL230 (ff05,06,07,08)
Aqawan B
0.4m A
SBIG (kb85)
0.4m B
FLI PL230 (ff09,10,11,12)
ASAS-SN
NRES Igloo

McDonald Observatory - Texas, USA

ELP

Dome A
1m
Sinistro (fa05)
NRES (nres-agu-ak16)
Dome B
1m
Sinistro (fa07)
NRES (nres-agu-ak18)
Aqawan A
0.4m A
None (sq31)
Telescope/Camera commissioning
0.4m B
FLI PL230 (ff17,18,19,20)
ASAS-SN
NRES Igloo

Haleakala Observatory - Maui, USA

OGG

Clamshell
2m
MuSCAT3 (mc03)
FLOYDS (en06)
0.4m A
FLI PL230 (ff01,02,03,04)
0.4m B
SBIG (kb27)
0.4m C
None (sq39)
Operational
Offline

Network-wide:

30 Jan 2023: The MARS database of ZTF alerts has been retired. MARS, launched in July 2018, was the first ZTF alert broker provided to the astronomical community. From its conception, MARS was intended as a bridge to other brokers, and now that they (ALeRCE, ANTARES, Lasair) are thriving, it was time for MARS to go.


Siding Spring

9-11 May 2023: The Spectral imager on the 2-meter telescope warmed up. The camera was taken offline for three nights.

6 Apr 2023: The 0m4a telescope, with the newly-installed kb26 SBIG camera, is available for science observations.

3 Apr 2023: Science observations with the FLOYDS spectrograph resume! Also, the kb26 camera has replaced kb23 on the 0m4a telescope.

2 Apr 2023: A new acquisition and guide camera (kb32) has been installed for the FLOYDS spectrograph. Camera testing is in-progress.

20 Mar 2023: The acquisition and guide camera on the FLOYDS spectrograph is producing extremely noisy images, making guiding impossible. FLOYDS is unavailable until a new camera is delivered and can be installed. The installation is currently planned for March 31.

26 Feb 2023: Moisture in the kb23 camera (0m4a telescope) has altered the field-of-view. Science data acquired on and after Feb 26 must be calibrated with flat fields acquired on and after this date.

20-23 Feb 2023: Problems with the electronics in the kb38 guide camera were preventing observations with the FLOYDS spectrograph. Only the Spectral imager was available on these nights.

10 Jan 2023: We've installed a new power supply for the time server, so we're getting our time stamps locally again.

30 Dec 2022: The site's time server went offline at 18:45 UT. Science observations continue, but time stamps are generated from a remote time source.

13-16 Dec 2022: A problem with the clamshell enclosure prevented observations with the 2-meter and 0.4-meter telescopes. Science observations resumed on the night of December 17.

15 Nov -3 Dec 2022: There was an intermittent problem (a failed solenoid) with the shutter on the FLOYDS spectrograph.

8 Nov 2022: In Dome B, the Safety PLC was replaced, and the door lock was repaired. The 1m telescope is available for science observations again.

25 Oct 2022: There's a fault on the Safety Controller of the Dome B telescope. This is another impediment to getting the telescope back online.

21 Oct 2022: Water has infiltrated the Dome B door switch, causing a breaker to trip. The 1m telescope in Dome B will be unavailable until either the site (and door switch) dries out, or the switch is replaced.

17 Oct 2022: The 2-meter telescope is available for science observations with both the Spectral camera and the FLOYDS spectrograph again.

7 Oct 2022: Maintenance on the 2-meter telescope has begun. The primary mirror will be recoated, and the Auxiliary Mechanisms Node will be replaced. This telescope is not available for science observations.

Wise

18-21 May 2023: A problem with the compressor in the cooling system kept the NRES unit offline.

21 Apr 2023: The NRES spectrograph is available for science observations again.

14 Apr 2023: All internet connections from outside of Israel through Tel Aviv University servers are blocked on this day.

23 Mar 2023: A leak has been detected in the NRES cryotiger cooler. NRES cannot operate until the leak is found and fixed.

18 Mar 2023: The NRES spectrograph has been taken offline for maintenance.

17 Feb 2023: A new acquisition camera (ak15) for NRES has been installed on the telescope. Test observations have concluded, and NRES is available for science observations again.

10 Feb 2023: The NRES unit is offline. There's a problem with the A/D conversion in the acquisition camera (ak14). The problem is under investigation.

29 Nov 2022: The mirrors of the telescope were washed today!

23 Oct 2022: A problem with the electronics crate supplying the acquisition camera is intermittently interrupting NRES observations.

17 Oct 2022: A communications/power problem with the acquisition camera forced NRES observations to be discontinued during the night.

Sutherland

29 May 2023: Lightning strikes have triggered a power outage at the SAAO. We have no communication to the site, and no observations are possible until observatory power is restored.

5 May 2023: Science validation observations with the 0m4a telescope (sq38 QHY camera) are underway.

4 May 2023: The NRES spectrograph is available for science observations!

2 May 2023: The 1m telescope in Dome C is available for science observations, but unfortunately, the telescope in Dome B is having difficulty tracking and has been taken offline.

1 May 2023: The 1m telescopes in Domes A & B are available for science observations.

21 Apr 2023: The air conditioning in the NRES igloo has been fixed, and the camera has been pumped down. Testing of the NRES spectrograph is underway.

18 Apr 2023: A team of LCO engineers has arrived on site!

21 Jan 2023: The 1m telescopes in Domes A & C are available for science observations again.

20 Jan 2023: The 1m telescope in Dome B is available for science observations again.

14 Jan 2023: The Archon controllers on the fa14 and fa06 Sinistro cameras have been replaced. Camera testing is underway.

13 Jan 2023: The loss of comms to the Dome B telescope was because multiple fibers had been chewed through. (Rats!) New fibers were installed. Communications to the Dome B telescope are restored.

12 Jan 2023: The mirrors on all three 1-meter telescopes were cleaned.

9 Jan 2023: A maintenance team from LCO HQ has arrived!

2 Dec 2022: The 1m telescope in Dome A (with the fa14 Sinistro camera) is available for science observations again.

29 Nov 2022: The SAAO staff replaced a controller on the secondary mirror of the 1m telescope in Dome C, and the telescope can now be focused again. The telescope has resumed science operations.

23 Nov 2022: The fa14 Sinistro camera has been installed on the 1m telescope in Dome A. Camera testing will begin after the Thanksgiving holiday.

17 Nov 2022: Control of the secondary mirror of the Dome C telescope has malfunctioned, and this telescope cannot be used for science observations until the problem is fixed. All three 1-meter telescopes at this site are now offline.

11 Nov 2022: We have lost communications to the Dome B telescope. Tests indicate that several optical fibers are broken.

7 Nov 2022: The fa06 Sinistro has cooled and been returned to service. Science observations with the Dome C telescope are underway.

1 Nov 2022: When it rains, it pours. Now the fa06 Sinistro camera (on the Dome C telescope) has warmed-up and can't be used for science observations.

27 Oct 2022: The fa16 Sinistro camera has warmed-up again. This camera (on the Dome A telescope) is not available for science observations.

2 Oct 2022: The fa16 camera has cooled and is available for science observations again.

Teide

8 May 2023: The 0m4a telescope isn't maintaining focus and has been taken offline.

20 Apr 2023: The sensors on the Dome B telescope's secondary mirror are working again. The Dome B telescope is now available for science observations.

9 Apr 2023: The secondary mirror on the Dome B telescope is stuck, preventing the telescope from focusing properly. This telescope is unavailable while the problem is investigated.

3 Apr 2023: The Dome A telescope is available for science observations again.

30 Mar 2023: An LCO engineer has arrived on site to service the cooling system on the Dome A Sinistro camera.

9-10 Mar 2023: The positioning of the secondary mirror on the Dome A telescope is unstable, intermittently creating out-of-focus images. This telescope was unavailable while the problem was under investigation.

2 Mar 2023: The Dome A telescope is available for science observations again.

27 Feb 2023: The Dome B telescope is available for science observations again.

9 Feb 2023: The telescope in Dome B is stuck on the pitch axis. The problem is under investigation.

6 Feb 2023: The Sinistro camera on the Dome A telescope has warmed-up and been taken offline.

30 Jan 2023: The encoders on the Dome A and Dome B telescopes have been upgraded.

28 Jan 2023: An LCO engineer has arrived at the site to undertake several maintenance tasks.

17 Nov 2022: The Teide Observatory staff cleaned the secondary mirror sensors, and the system is behaving itself. The telescope is available for science observations again.

27 Oct 2022: There's a problem communicating with the secondary mirror of the Dome B telescope, affecting the telescope's focus. This telescope is offline.

8 Oct 2022: The kb24 camera on the 0m4b telescope has been fixed, and it's now available for science observations.

Cerro Tololo

19 Apr 2023: A coupling on one of the secondary mirror controllers of the 0m4a telescope in Aqawan A has been replaced. The AqwA:0m4a telescope is available for science observations again.

29 Mar 2023: The 0m4a telescope in Aqawan A won't focus. This telescope is offline while the problem is investigated.

14 Feb 2023: The fa03 Sinistro camera on the Dome C telescope has cooled and is available for science observations.

9 Feb 2023: The fa03 Sinistro camera on the Dome C telescope has warmed-up and been taken offline. Observations with NRES on this telescope continue.

4 Jan 2023: The 0.4m telescope in Aqawan is available for science observations again.

22 Dec 2022: Problems with the focus mechanism on the 0.4m telescope in Aqawan B have suspended observations with this telescope.

30 Nov 2022: The CTIO staff have repaired the focus mechanism on the 0.4m telescope in Aqawan A (camera kb98), and the telescope is available for science observations again.

14 Nov 2022: The 0.4m telescope (with the kb85 SBIG camera) in Aqawan B is available for science observations again. Unfortunately, there's a problem with the focus mechanism on the 0.4m telescope in Aqawan A.

9 Nov 2022: The 1m telescope in Dome A (with the fa15 Sinistro camera) is now available for science observations.

4 Nov 2022: The 1m telescope in Dome C (for the fa03 Sinistro image and the NRES spectrograph) is available for science observations again.

3 Nov 2022: The 1m telescope in Dome B (with the fa04 Sinistro camera) is available for science observations again.

23 Oct 2022: The fa15 Sinistro camera, on the Dome A telescope, has warmed-up. The maintenance work at LCO's site has shifted into high gear. Only the 0m4a telescope in Aqawan A will be available for science observations for approximately a week.

21 Oct 2022: Maintenance on the NRES spectrograph and the Dome C telescope has begun. These assets will be unavailable for the next several days.

20 Oct 2022: An LCO engineering team has arrived at site.

14 Oct 2022: The NRES spectrograph is available for science observations with the Dome C telescope providing the signal.

11 Oct 2022: Now the Dome C telescope has recovered from the power outage and is available for science observations. NRES is still offline.

9 Oct 2022: The Dome A telescope has recovered from the power outage and is available for science observations again tonight.

8 Oct 2022: Only the 0m4a telescope in Aqawan A is available for science observations tonight.

7 Oct 2022: Another power outage at CTIO. Observations during the night were compromised.

6 Oct 2022: There was a power outage at Cerro Tololo during the night. The power was later restored, but the NRES spectrograph will take a few days to reach temperature.

McDonald

29 May 2023: An LCO engineer has arrived on site to improve the collimation of the 0m4a telescope. The sq31 camera replaced sq32 on that telescope.

10 Mar 2023: The ELP-NRES unit is available for science observations again.

3 Mar 2023: The ELP-NRES unit has been taken offline because of temperature instabilities in the igloo following the site-wide power outages.

1 Mar 2023: A QHY600 camera (sq32) has been installed in place of the kb55 SBIG camera. Testing is in progress.

28 Feb 2023: A DeltaRho 350 telescope has been installed as 0m4a.

26 Feb 2023: Because of the nearby Solar Park Wildfire, McDonald Observatory has no power or internet. Early on Feb 27, the Fort Davis Fire Dept reported that there was no imminent danger to Observatory structures.

20 Jan 2023: The 1m telescope in Dome A is available for science observations again.

18 Jan 2023: The ef12 guide camera on the Dome A telescope has been replaced by ef08. Testing is underway.

17 Jan 2023: The ak07 NRES acquisition camera on the Dome B telescope has been replaced by ak18. Testing is underway.

16 Jan 2023: The aqawan shutter has been repaired. The 0m4a telescope will be checked-out before science observations resume.

15 Jan 2023: A maintenance team from LCO HQ has arrived at McDonald Observatory.

23 Dec 2022: There's a mechanical problem with one of the shutters of the aqawan enclosure. Observations with the 0.4m telescope and the ASAS-SN telescope are suspended until the aqawan can be safely opened and closed. The mechanical problem may have been induced by the sudden drop to extremely low (-20 C) temperatures.

Haleakala

5 May 2023: This is the first night that the Haleakala site is operating with the newly-installed weather system.

24 Apr 2023: A new Vaisala WXT536 weather sensor has been delivered to the site, and testing has begin. (This will replace the weather sensor that was struck by lightning in January.)

7 Apr 2023: A reliable time server has been restored! The timing glitches have ceased, and the tracking errors eliminated.

4 Apr 2023: The kb41 guide camera was re-installed on the 2-meter telescope. Science observations with FLOYDS resume.

29 Mar 2023: Problems with the filter wheel on the kb31 guider are keeping FLOYDS offline.

28 Mar 2023: The kb31 camera was installed in place of kb41 on FLOYDS. Testing of the camera is in progress (when weather permits).

17 Mar 2023: The kb41 acquisition and guide camera for FLOYDS is intermittently getting stuck in readout mode. A replacement camera was sent on March 21.

16 Mar 2023: Timing glitches on the altitude axis of the 2-meter telescope are intermittently causing observations to fail when tracking errors exceed their tolerance. This problem will be eliminated when a new time server is installed in April.

4 Mar 2023: Another power outage on Haleakala.

22-24 Feb 2023: There was a problem with the oil system on the 2m telescope. (Bad weather was a problem, too!) The leaked oil has been cleaned-up, and the telescope can resume science observations when the weather clears.

8 Feb 2023: We have reconfigured our site software to draw weather data from various sources. Science observations will only be made when our Telescope Operations team has checked weather forecasts and inspected the weather data.

30 Jan 2023: The lightning blasted our weather station, as well as other systems on site. A damage assessment is underway.

28 Jan 2023: A lightning storm(!) at Haleakala Observatory has knocked out our ability to continue observations. Fortunately(?!?), conditions at the site are terrible for astronomy (rain, snow).

23 Jan 2023: The 0m4b telescope (with the kb27 camera) is available for science observations again.

19 Jan 2023: Science observations resume! Because the site has been offline for a month, final images/spectra will be reprocessed when new calibration observations are acquired.

16 Jan 2023: Outrage! Frustration! Communications to a router dropped when an adjacent U. Hawaii building lost power.

15 Jan 2023: Manual inspection of the clamshell enclosure and the 2-meter telescope is complete. Instrument testing is next.

13 Jan 2023: The fiber to the observatory was patched, restoring communications.

27 Dec 2022: Power to Haleakala Observatory has been restored. However, our fiber connection has been severed by a landslide on the west side of the summit. Observations are still not possible.

18 Dec 2022: A storm has knocked out power to Haleakala Observatory. No observations are possible until power is restored.

4 Dec 2022: A south wind blew smoke from the Mauna Loa eruption toward Maui, and the observatory remained closed.

11 Nov 2022: A new acquisition camera has been installed for FLOYDS. Science observations have resumed.

2 Nov 2022: There's a problem communicating with the FLOYDS acquisition camera. FLOYDS will be offline until the problem is fixed.